# Distribution-Aligned Decoding for Efficient LLM Task Adaptation

### v3

@@ Abstract
Adapting billion-parameter language models to a downstream task is still costly, even with parameter-efficient fine-tuning (PEFT). We re-cast task adaptation as output-distribution alignment: the objective is to steer the output distribution toward the task distribution directly during decoding rather than indirectly through weight updates. Building on this view, we introduce Steering Vector Decoding (SVDecode), a lightweight, PEFT-compatible, and theoretically grounded method. We start with a short warm-start fine-tune and extract a task-aware steering vector from the Kullback-Leibler (KL) divergence gradient between the output distribution of the warm-started and pre-trained models. This steering vector is then used to guide the decoding process to steer the model’s output distribution towards the task distribution. We theoretically prove that SVDecode is first-order equivalent to the gradient step of full fine-tuning and derive a globally optimal solution for the strength of the steering vector. Across three tasks and nine benchmarks, SVDecode paired with four standard PEFT methods improves multiple-choice accuracy by up to 5 percentage points and open-ended truthfulness by 2 percentage points, with similar gains (1-2 percentage points) on commonsense datasets without adding trainable parameters beyond the PEFT adapter. SVDecode thus offers a lightweight, theoretically grounded path to stronger task adaptation for large language models.
@@ 1 Introdcution
Large language models (LLMs) [1, 2, 3, 4] are pivotal in AI, marking early steps towards artificial general intelligence (AGI). They excel in tasks like language understanding, generation, and translation, transforming natural language processing with their grasp of context and human intent. Models like DeepSeek-R1 [4] and OpenAI o1 [5] demonstrate strong reasoning and multimodal capabilities, respectively. Specialized models such as EmoLLMs [6], LMDrive [7], and AnomalyGPT [8] address specific downstream tasks like affective instructions, autonomous driving, and anomaly detection. Despite their capabilities, LLMs are resource-intensive, often requiring hundreds of millions to billions of parameters. For instance, training a LLaMA-7B model demands at least 58 GB of memory [9], which is beyond the capacity of consumer-grade hardware like the NVIDIA RTX 4090 with 24GB, limiting their broader applications.
To tackle this challenge, parameter-efficient fine-tuning (PEFT) [10, 11, 12, 13, 14, 15] has emerged as a key area of progress in modifying LLMs with minimal computational and GPU memory demands. This approach focuses on updating a few trainable parameters to significantly reduce the memory footprint while enhancing the performance on downstream tasks. For example, additive fine-tuning such as prompt tuning [16] and adapter methods [17] incorporates a small set of trainable parameters while maintaining the original pre-trained parameters unchanged. Selective fine-tuning such as Diff Pruning [18] chooses a subset of the model’s existing parameters to undergo updates during training. Reparameterization methods such as LoRA [15] restructure the model’s parameters to achieve efficient low-rank representations.
@@ 2 Rethinking LLM Task Adaptation from the Perspective of Output Distribution Alignment
Distributional Interpretation. From the output distribution perspective, fine-tuning reshapes the model’s belief $P_{\theta}(y|x)$ over the output space to align more closely with the true task-specific behavior. The output distribution $P_{\theta}(y|x)$ resides on the probability simplex $\Delta^{|\mathcal{Y}|-1}$ [19], and fine-tuning can be seen as shifting the model’s position on this simplex toward the optimal region defined by the task. Minimizing the KL divergence from the empirical distribution emphasizes increasing the probability mass on the correct label without penalizing overconfidence in incorrect predictions. This yields a learning dynamic that is both efficient and focused.
@@ 4.1 Experimental Setup
    Multiple-Choice Tasks. For multiple-choice and open-ended generation tasks, we evaluate on the TruthfulQA dataset [21], which is a benchmark designed to measure a model’s tendency to generate truthful answers to questions. We consider three metrics in this task: MC1, MC2, and MC3. The detailed definitions of these metrics are shown in Appendix E.3.
@@
    Open-Ended Generation Tasks. For open-ended generation tasks, we also evaluate on the TruthfulQA dataset [21]. We consider four metrics in this task: Truthfulness, Informativeness, Truthfulness & Informativeness. The detailed definitions of these metrics are shown in Appendix E.3.
@@
    Commonsense Reasoning Tasks. For commonsense reasoning tasks, we leverage eight datasets including BoolQ [22], PIQA [23], SIQA [24], HellaSwag [25], WinoGrande [26], ARC-easy [27], ARC-challenge [27] and OBQA [28], and we leverage accuracy as the metric. The implementation details are shown in Appendix E.5.
Base Models and PEFT Methods. We consider four latest pre-trained LLMs: Qwen2.5-1.5B [2], Qwen2.5-7B [2], LLaMA3-8B [1], and LLaMA3.1-8B [1] as the base models. In addition, we leverage four PEFT methods to incorporate our method: LoRA [29], P-Tuning v2 [30], Prompt Tuning [16], and IA3 [31]. We eloborate the implementation details in Appendix E.
@@ 6 Acknowledgement
The research work described in this paper was conducted in the JC STEM Lab of Smart City funded by The Hong Kong Jockey Club Charities Trust under Contract 2023-0108. The work was supported in part by the Hong Kong SAR Government under the Global STEM Professorship and Research Talent Hub. The work of Senkang Hu was supported in part by the Hong Kong Innovation and Technology Commission under InnoHK Project CIMDA.
@@ A.1 Parameter-Efficient Fine-Tuning
As language models continue to increase in scale, traditional full fine-tuning approaches have become increasingly resource-intensive. Parameter-efficient fine-tuning (PEFT) emerges as a practical alternative to address these computational constraints [11, 12, 13, 14, 15]. According to [32], PEFT techniques generally fall into three distinct categories: 1) Additive Fine-Tuning, which incorporates a limited set of trainable parameters while maintaining the original pre-trained parameters unchanged. Notable examples in this category include Adapter modules [33] and Prompt Tuning, which integrates learnable soft prompts into the input. Though effective, these approaches typically introduce additional computational demands during inference. 2) Selective Fine-Tuning, which focuses on updating only a carefully chosen subset of the model’s existing parameters [18, 34, 35]. This strategy employs a binary mask $\mathcal{M}$ to selectively determine which parameters undergo updates during training, thereby avoiding the introduction of new parameters. 3) Reparameterization, which restructures the model’s parameters to achieve efficient low-rank representations [15, 36, 14]. A prominent example is LoRA [15], which factorizes weight updates into products of smaller matrices, facilitating compact storage of task-specific adaptations. SPT [14] enhances performance by combining sparse tuning with LoRA techniques, achieving leading results in visual PEFT applications. Nevertheless, many existing approaches still lack sufficient task-awareness in their parameter selection mechanisms.
@@ A.2 LLM Task Adaptation
Researchers have developed diverse approaches to adapt pre-trained LLMs for downstream tasks by investigating optimal configurations of pre-training data, model architecture, and weight selection. For instance, Cui et al. [37] leverage Earth Mover’s Distance to select the top $K$ most relevant classes from the source domain for task-specific pre-training. Yoon et al. [38] apply reinforcement learning techniques to determine appropriate weights for source domain classes. Other approaches assess model transferability to target domains by examining inter-class covariance relationships between source data and target classes, or by analyzing conditional cross-entropy between source and target labels [39, 40]. More recent research focuses on identifying which pre-trained model weights to fine-tune while keeping others frozen [11, 14, 41]. Zhang et al. [11] introduced gradient-based parameter selection (GPS), a method that uses gradients to identify optimal tunable weights. Fu et al. [41] developed a second-order approximation method (SAM) that approximates the Hessian matrix in the loss function’s second-order Taylor expansion to enhance weight selection precision.
@@ A.3 LLM Decoding Strategies
Decoding methods play a pivotal role in shaping the output characteristics of large language models, determining both the fluency and creativity of generated text. The most basic Greedy Search approach has been observed to frequently fall into repetitive patterns due to its myopic selection strategy. More advanced methods like Beam Search [42] have addressed this limitation by exploring multiple potential sequences simultaneously, though at increased computational cost. Compared to greedy search’s tendency for repetitive outputs and beam search’s limited diversity, Top-k sampling [43] introduces diversity by sampling from a fixed-size set of most probable tokens, while Top-p sampling [44] further improves adaptability by dynamically adjusting the candidate set based on probability mass. Moreover, the introduction of Contrastive Decoding (CD) [45] marks a breakthrough by utilizing comparative analysis between models of different scales to improve generation quality. Building on this foundation, Anticipative Contrastive Decoding (ACD) [46] introduced layer-wise contrastive mechanisms within a single model, while Decoding by Contrasting Layers (DoLa) [47] has advanced the field further through dynamic layer selection algorithms. Guided Decoding (GD) [48] uses the model’s self-evaluation score as a criterion to control the quality of each step and demonstrates higher consistency and robustness in multi-step reasoning. These decoding approaches, however, predominantly emphasize optimizing pre-trained LLM performance without accounting for model transformations that occur during fine-tuning. As a result, they inadequately leverage the task-specific adaptations acquired through fine-tuning processes, which often leads to suboptimal performance gains or even degradation when implemented with fine-tuned LLMs on downstream applications. On the contrary, our method focuses on LLM task adaptation via decoding with task-aware steered vectors, thereby enhancing the performance of LLMs in downstream tasks.
@@ E.3 Evaluation Metrics
To evaluate the performance of our method on open-ended generation tasks, we consider Truthfulness, Informativeness, and Truthfulness & Informativeness. Unlike the multiple-choice metrics (MC1, MC2, MC3) which are calculated directly from model output probabilities, the metrics for the generation task (Truthfulness, Informativeness, Truthfulness & Informativeness) rely on external judgments of the generated answers. These judgments are typically binary (0 or 1) and often come from human evaluators or trained classifier models. In our experiments, we use DeepSeek-V3-0324 [3] as the external judge.
@@ E.4 Implementation Details on Multiple-Choice Tasks and Open-Ended Generation Tasks
In order to evaluate the performance of our method on multiple-choice tasks and open-ended generation tasks, we leverage TruthfulQA dataset [21] as the training and evaluation dataset. For multiple-choice tasks, we use the MC1, MC2, and MC3 metrics. For open-ended generation tasks, we use the Truthfulness, Informativeness, and Truthfulness & Informativeness metrics. The details of the metrics are shown in Appendix E.3. In addition, we leverage DeepSeek-V3-0324 [3] as the external judge. The details are shown in Appendix E.6.
@@ E.5 Implementation Details on Commonsense Reasoning Tasks
In order to evaluate the performance of our method on commonsense reasoning tasks, we leverage eight datasets including BoolQ [22], PIQA [23], SIQA [24], HellaSwag [25], WinoGrande [26], ARC-easy [27], ARC-challenge [27] and OBQA [28], using accuracy as the metric. Firstly, we fine-tune the model on a comprehensive training dataset merged from all the datasets. Then, we evaluate the method on each task’s test set. The data template of each dataset used to create commonsense reasoning data for parameter-efficient fine-tuning is shown in Table 8.
We fine-tune three models including Qwen2.5-7B [2], LLaMA3-8B [1], and LLaMA3.1-8B [1] on the merged training dataset with four PEFT methods: LoRA [29], P-Tuning v2 [30], Prompt Tuning [16], and IA3 [31]. The hyperparameters are summarized in Table 7.
@@ E.6 Details about DeepSeek-V3-0324 Evaluation
To evaluate the performance of our method on open-ended generation tasks, traditional approaches are to use human evaluators or train classifier models to judge the quality of the generated answers. However, this method is inefficient and costly. LLMs with strong reasoning capabilities, such as GPT-4 and DeepSeek-R1/V3, have been proven to be an alternative to human evaluation in many cases with stable performance over different prompts and instructions [49, 50]. Here, we use DeepSeek-V3-0324 [3] as the external judge. The prompt used for evaluation is shown in Table 9. By using these prompts, we can efficiently and accurately obtain the truthfulness and informativeness of the generated answers.

### v5

# arXiv:2509.15888v5[cs.CL] 28 Feb 2026

@@ Abstract
Adapting billion-parameter language models to a downstream task is still costly, even with parameter-efficient fine-tuning (PEFT). We re-cast task adaptation as output-distribution alignment: the objective is to steer the output distribution toward the task distribution directly during decoding rather than indirectly through weight updates. Building on this view, we introduce Steering Vector Decoding (SVDecode), a lightweight, PEFT-compatible, and theoretically grounded method. We start with a short warm-start fine-tune and extract a task-aware steering vector from the Kullback-Leibler (KL) divergence gradient between the output distribution of the warm-started and pre-trained models. This steering vector is then used to guide the decoding process to steer the model’s output distribution towards the task distribution. We theoretically prove that SVDecode is first-order equivalent to the gradient step of full fine-tuning and derive a globally optimal solution for the strength of the steering vector. Across three tasks and nine benchmarks, SVDecode paired with four standard PEFT methods improves multiple-choice accuracy by up to 5 points and open-ended truthfulness by 2 points, with similar gains (1-2 points) on commonsense datasets without adding trainable parameters beyond the PEFT adapter. SVDecode thus offers a lightweight, theoretically grounded path to stronger task adaptation for large language models. Code is available at https://github.com/dl-m9/SVDecode.
@@ 1 Introdcution
Large language models (LLMs) [llama3modelcard, qwen2, deepseekai2025deepseekv3technicalreport, deepseekai2025deepseekr1incentivizingreasoningcapability] are pivotal in AI, marking early steps towards artificial general intelligence (AGI). They excel in tasks like language understanding, generation, and translation, transforming natural language processing with their grasp of context and human intent. Models like DeepSeek-R1 [deepseekai2025deepseekr1incentivizingreasoningcapability] and OpenAI o1 [openai2024openaio1card] demonstrate strong reasoning and multimodal capabilities, respectively. Specialized models such as EmoLLMs [10.1145/3637528.3671552], LMDrive [10657019], and AnomalyGPT [10.1609/aaai.v38i3.27963] address specific downstream tasks like affective instructions, autonomous driving, and anomaly detection. Despite their capabilities, LLMs are resource-intensive, often requiring hundreds of millions to billions of parameters. For instance, training a LLaMA-7B model demands at least 58 GB of memory [10.5555/3692070.3694598], which is beyond the capacity of consumer-grade hardware like the NVIDIA RTX 4090 with 24GB, limiting their broader applications.
To tackle this challenge, parameter-efficient fine-tuning (PEFT) [hu2025taskawareparameterefficientfinetuninglarge, 9, 7, 1, 4, hu2022lora] has emerged as a key area of progress in modifying LLMs with minimal computational and GPU memory demands. This approach focuses on updating a few trainable parameters to significantly reduce the memory footprint while enhancing the performance on downstream tasks. For example, additive fine-tuning such as prompt tuning [lester-etal-2021-power] and adapter methods [6] incorporates a small set of trainable parameters while maintaining the original pre-trained parameters unchanged. Selective fine-tuning such as Diff Pruning [guo-etal-2021-parameter] chooses a subset of the model’s existing parameters to undergo updates during training. Reparameterization methods such as LoRA [hu2022lora] restructure the model’s parameters to achieve efficient low-rank representations.
@@ 2 Rethinking LLM Task Adaptation from the Perspective of Output Distribution Alignment
Distributional Interpretation. From the output distribution perspective, fine-tuning reshapes the model’s belief $P_{\theta}(y|x)$ over the output space to align more closely with the true task-specific behavior. The output distribution $P_{\theta}(y|x)$ resides on the probability simplex $\Delta^{|\mathcal{Y}|-1}$ [borgwardt2012simplex], and fine-tuning can be seen as shifting the model’s position on this simplex toward the optimal region defined by the task. Minimizing the KL divergence from the empirical distribution emphasizes increasing the probability mass on the correct label without penalizing overconfidence in incorrect predictions. This yields a learning dynamic that is both efficient and focused.
@@ 4.1 Experimental Setup
    Multiple-Choice Tasks. For multiple-choice and open-ended generation tasks, we evaluate on the TruthfulQA dataset [lin-etal-2022-truthfulqa], which is a benchmark designed to measure a model’s tendency to generate truthful answers to questions. We consider three metrics in this task: MC1, MC2, and MC3. The detailed definitions of these metrics are shown in Appendix E.3.
@@
    Open-Ended Generation Tasks. For open-ended generation tasks, we also evaluate on the TruthfulQA dataset [lin-etal-2022-truthfulqa]. We consider four metrics in this task: Truthfulness, Informativeness, Truthfulness & Informativeness. The detailed definitions of these metrics are shown in Appendix E.3.
@@
    Commonsense Reasoning Tasks. For commonsense reasoning tasks, we leverage eight datasets including BoolQ [BoolQ], PIQA [PIQA], SIQA [SIQA], HellaSwag [HellaSwag], WinoGrande [WinoGrande], ARC-easy [ARC], ARC-challenge [ARC] and OBQA [OBQA], and we leverage accuracy as the metric. The implementation details are shown in Appendix E.5.
Base Models and PEFT Methods. We consider four latest pre-trained LLMs: Qwen2.5-1.5B [qwen2], Qwen2.5-7B [qwen2], LLaMA3-8B [llama3modelcard], and LLaMA3.1-8B [llama3modelcard] as the base models. In addition, we leverage four PEFT methods to incorporate our method: LoRA [5], P-Tuning v2 [liu-etal-2022-p], Prompt Tuning [lester-etal-2021-power], and IA3 [10.5555/3600270.3600412]. We eloborate the implementation details in Appendix E.
@@ 6 Acknowledgement
The research work described in this paper was conducted in the JC STEM Lab of Smart City funded by The Hong Kong Jockey Club Charities Trust under Contract 2023-0108. The work was supported in part by the Hong Kong SAR Government under the Global STEM Professorship and Research Talent Hub. The work of S. Hu was supported in part by the Hong Kong Innovation and Technology Commission under InnoHK Project CIMDA.
@@ A.1 Parameter-Efficient Fine-Tuning
As language models continue to increase in scale, traditional full fine-tuning approaches have become increasingly resource-intensive. Parameter-efficient fine-tuning (PEFT) emerges as a practical alternative to address these computational constraints [9, 7, 1, 4, hu2022lora]. According to [3], PEFT techniques generally fall into three distinct categories: 1) Additive Fine-Tuning, which incorporates a limited set of trainable parameters while maintaining the original pre-trained parameters unchanged. Notable examples in this category include Adapter modules [pmlr-v97-houlsby19a] and Prompt Tuning, which integrates learnable soft prompts into the input. Though effective, these approaches typically introduce additional computational demands during inference. 2) Selective Fine-Tuning, which focuses on updating only a carefully chosen subset of the model’s existing parameters [guo-etal-2021-parameter, das-etal-2023-unified, liao-etal-2023-parameter]. This strategy employs a binary mask $\mathcal{M}$ to selectively determine which parameters undergo updates during training, thereby avoiding the introduction of new parameters. 3) Reparameterization, which restructures the model’s parameters to achieve efficient low-rank representations [hu2022lora, aghajanyan-etal-2021-intrinsic, 4]. A prominent example is LoRA [hu2022lora], which factorizes weight updates into products of smaller matrices, facilitating compact storage of task-specific adaptations. SPT [4] enhances performance by combining sparse tuning with LoRA techniques, achieving leading results in visual PEFT applications. Nevertheless, many existing approaches still lack sufficient task-awareness in their parameter selection mechanisms.
@@ A.2 LLM Task Adaptation
Researchers have developed diverse approaches to adapt pre-trained LLMs for downstream tasks [10779389, Hu_Tao_Xu_Deng_Chen_Fang_Kwong_2025, shao2023lmdrive, tao2025directedcpdirectedcollaborativeperception, 11020620, 10976336, yang2024emollmmultimodalemotionalunderstanding, hu2024agentscodriverlargelanguagemodel, hu2025cpguardnewparadigmmalicious, hu2025cpuniguardunifiedprobabilityagnosticadaptive, gu2023anomalyagpt] by investigating optimal configurations of pre-training data, model architecture, and weight selection. For instance, Cui et al. [8578530] leverage Earth Mover’s Distance to select the top $K$ most relevant classes from the source domain for task-specific pre-training. Yoon et al. [pmlr-v119-yoon20a] apply reinforcement learning techniques to determine appropriate weights for source domain classes. Other approaches assess model transferability to target domains by examining inter-class covariance relationships between source data and target classes, or by analyzing conditional cross-entropy between source and target labels [9009545, pmlr-v119-nguyen20b]. More recent research focuses on identifying which pre-trained model weights to fine-tune while keeping others frozen [9, 4, 2]. Zhang et al. [9] introduced gradient-based parameter selection (GPS), a method that uses gradients to identify optimal tunable weights. Fu et al. [2] developed a second-order approximation method (SAM) that approximates the Hessian matrix in the loss function’s second-order Taylor expansion to enhance weight selection precision.
@@ A.3 LLM Decoding Strategies
Decoding methods play a pivotal role in shaping the output characteristics of large language models, determining both the fluency and creativity of generated text. The most basic Greedy Search approach has been observed to frequently fall into repetitive patterns due to its myopic selection strategy. More advanced methods like Beam Search [sutskever2014sequence] have addressed this limitation by exploring multiple potential sequences simultaneously, though at increased computational cost. Compared to greedy search’s tendency for repetitive outputs and beam search’s limited diversity, Top-k sampling [topk] introduces diversity by sampling from a fixed-size set of most probable tokens, while Top-p sampling [topp] further improves adaptability by dynamically adjusting the candidate set based on probability mass. Moreover, the introduction of Contrastive Decoding (CD) [cd_decode] marks a breakthrough by utilizing comparative analysis between models of different scales to improve generation quality. Building on this foundation, Anticipative Contrastive Decoding (ACD) [acd_decode] introduced layer-wise contrastive mechanisms within a single model, while Decoding by Contrasting Layers (DoLa) [dola_decode] has advanced the field further through dynamic layer selection algorithms. Guided Decoding (GD) [gd_decode] uses the model’s self-evaluation score as a criterion to control the quality of each step and demonstrates higher consistency and robustness in multi-step reasoning. These decoding approaches, however, predominantly emphasize optimizing pre-trained LLM performance without accounting for model transformations that occur during fine-tuning. As a result, they inadequately leverage the task-specific adaptations acquired through fine-tuning processes, which often leads to suboptimal performance gains or even degradation when implemented with fine-tuned LLMs on downstream applications. On the contrary, our method focuses on LLM task adaptation via decoding with task-aware steered vectors, thereby enhancing the performance of LLMs in downstream tasks.
@@ E.3 Evaluation Metrics
To evaluate the performance of our method on open-ended generation tasks, we consider Truthfulness, Informativeness, and Truthfulness & Informativeness. Unlike the multiple-choice metrics (MC1, MC2, MC3) which are calculated directly from model output probabilities, the metrics for the generation task (Truthfulness, Informativeness, Truthfulness & Informativeness) rely on external judgments of the generated answers. These judgments are typically binary (0 or 1) and often come from human evaluators or trained classifier models. In our experiments, we use DeepSeek-V3-0324 [deepseekai2025deepseekv3technicalreport] as the external judge.
@@ E.4 Implementation Details on Multiple-Choice Tasks and Open-Ended Generation Tasks
In order to evaluate the performance of our method on multiple-choice tasks and open-ended generation tasks, we leverage TruthfulQA dataset [lin-etal-2022-truthfulqa] as the training and evaluation dataset. For multiple-choice tasks, we use the MC1, MC2, and MC3 metrics. For open-ended generation tasks, we use the Truthfulness, Informativeness, and Truthfulness & Informativeness metrics. The details of the metrics are shown in Appendix E.3. In addition, we leverage DeepSeek-V3-0324 [deepseekai2025deepseekv3technicalreport] as the external judge. The details are shown in Appendix E.6.
@@ E.5 Implementation Details on Commonsense Reasoning Tasks
In order to evaluate the performance of our method on commonsense reasoning tasks, we leverage eight datasets including BoolQ [BoolQ], PIQA [PIQA], SIQA [SIQA], HellaSwag [HellaSwag], WinoGrande [WinoGrande], ARC-easy [ARC], ARC-challenge [ARC] and OBQA [OBQA], using accuracy as the metric. Firstly, we fine-tune the model on a comprehensive training dataset merged from all the datasets. Then, we evaluate the method on each task’s test set. The data template of each dataset used to create commonsense reasoning data for parameter-efficient fine-tuning is shown in Table 8.
We fine-tune three models including Qwen2.5-7B [qwen2], LLaMA3-8B [llama3modelcard], and LLaMA3.1-8B [llama3modelcard] on the merged training dataset with four PEFT methods: LoRA [5], P-Tuning v2 [liu-etal-2022-p], Prompt Tuning [lester-etal-2021-power], and IA3 [10.5555/3600270.3600412]. The hyperparameters are summarized in Table 7.
@@ E.6 Details about DeepSeek-V3-0324 Evaluation
To evaluate the performance of our method on open-ended generation tasks, traditional approaches are to use human evaluators or train classifier models to judge the quality of the generated answers. However, this method is inefficient and costly. LLMs with strong reasoning capabilities, such as GPT-4 and DeepSeek-R1/V3, have been proven to be an alternative to human evaluation in many cases with stable performance over different prompts and instructions [chiang-lee-2023-large, chiang-lee-2023-closer]. Here, we use DeepSeek-V3-0324 [deepseekai2025deepseekv3technicalreport] as the external judge. The prompt used for evaluation is shown in Table 9. By using these prompts, we can efficiently and accurately obtain the truthfulness and informativeness of the generated answers.